OpenStudy (anonymous):

Find the remainder when f(x) is divided by (x - k) f(x) = 3x3 - 4x2 - 3x + 14; k= 3 A. 50 B. 68 C. -12 D. 112

OpenStudy (campbell_st):

you can simply find f(3) and that will give the remainder. the process is using the remainder theorem.

OpenStudy (anonymous):

I dont know what the remained theorem is..

OpenStudy (campbell_st):

ok but you can find f(3) can't you..?

OpenStudy (anonymous):

how would i do that?

OpenStudy (campbell_st):

the answer will be the remainder...

OpenStudy (campbell_st):

just like this \[f(3) = 3\times(3)^3 - 4 \times (3)^2 - 3 \times (3) + 14\]

OpenStudy (campbell_st):

the answer will be the remainder...

OpenStudy (anonymous):

ohh so just replacing the xs for 3's

OpenStudy (campbell_st):

thats it... then evaluating

OpenStudy (anonymous):

i got 2 ?

OpenStudy (campbell_st):

maybe you want to check that... my answer is one of the choices

OpenStudy (campbell_st):

\[f(3) = 3 \times 27 - 4 \times 9 - 3 \times 3 + 14\]

OpenStudy (anonymous):

hold on yeah nvm

OpenStudy (anonymous):

i got 50? answer A

OpenStudy (campbell_st):

that correct... well done

OpenStudy (anonymous):

thank you again.. by anychance do u no how to do this.. Use synthetic division to determine whether the number k is an upper or lower bound (as specified) for the real zeros of the function f. k = -1; f(x) = 4x3 - 2x2 + 2x + 4; Lower bound? Yes No

OpenStudy (anonymous):

i got yes but im not sure?

OpenStudy (anonymous):

if you dnt no its ok.. :)

OpenStudy (campbell_st):

I've not done synthetic division...

OpenStudy (anonymous):

oh ok thnku nyways

OpenStudy (anonymous):

for anyone needing the answer to this: it's 50 :)